### v1

# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

Abstract
Let $S_{n}$ and $S_{n,k}$ be, respectively, the number of subsets and $k$-subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ such that no two subset elements differ by an element of the set $\mathcal{Q}$. We prove a bijection between such $k$-subsets when $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ and permutations $\pi$ of $\mathbb{N}_{n+jm}$ with $k$ excedances satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$. We also identify a bijection between another class of restricted permutation and the cases $\mathcal{Q}=\{1,q\}$. This bijection allows us to prove a conjectured recursion relation for the number of such permutations which corresponds to the case $\mathcal{Q}=\{1,4\}$. We also obtain recursion relations for $S_{n}$ and $S_{n,k}$ in the case $\mathcal{Q}=\{1,5\}$ by first obtaining related recursion relations for the numbers of closed walks of a given length on a particular class of directed pseudograph. We give some classes of $\mathcal{Q}$ for which $S_{n}$ is also the number of compositions of $n+q$ into a given set of allowed parts, where $q$ is the largest element of $\mathcal{Q}$. A bijection between the $k$-subsets for any $\mathcal{Q}$ and bit strings is also noted. Aided by this, an efficient algorithm for finding $S_{n}$ and $S_{n,k}$ is given. We also prove a bijection between $k$-subsets for a class of $\mathcal{Q}$ and the set representations of size $k$ of equivalence classes for the occurrence of a given length-($q+1$) subword within bit strings. We then formulate a straightforward procedure for obtaining the generating function for the number of such equivalence classes.
1 Introduction
Combinations without specified separations (or restricted combinations) refer to the subsets of $\mathbb{N}_{n}=\{1,\ldots,n\}$ with the property that no two elements of a subset differ by an element of a set $\mathcal{Q}$, the elements of which are independent of $n$. We denote the total number of such subsets and the number of $k$-subsets (i.e., subsets of size $k$) by $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$, respectively. We drop the superscripts when it is clear which set of disallowed differences they apply to. For any $\mathcal{Q}$, one always takes $S^{\mathcal{Q}}_{n<0}=S^{\mathcal{Q}}_{n,k<0}=S^{\mathcal{Q}}_{n<k,k}=0$. When there are no restrictions, we just have the number of ordinary combinations and so $S^{\{\}}_{n}=2^{n}$ (A000079 in the OEIS [20]) and $S^{\{\}}_{n,k}=\tbinom{n}{k}$ (A007318). It appears to be well-known that $S^{\{1,\ldots,q\}}_{n}=S^{\{1,\ldots,q\}}_{n-1}+S^{\{1,\ldots,q\}}_{n-q-1}+\delta_{n+q,0}$, where $\delta_{i,j}$ is 1 if $i=j$ and 0 otherwise, and $S^{\{1,\ldots,q\}}_{n,k}=\tbinom{n+q(1-k)}{k}$. These reduce to the classic results $S^{\{1\}}_{n}=F_{n+2}$, where $F_{n}$ is the $n$-th Fibonacci number (A000045, $F_{n}=F_{n-1}+F_{n-2}+\delta_{n,1}$, $F_{n<1}=0$), and $S^{\{1\}}_{n,k}=\tbinom{n+1-k}{k}$ (A011973) [12]. Expressions for $S_{n}^{\mathcal{Q}}$ and $S^{\mathcal{Q}}_{n,k}$ have also been obtained when $\mathcal{Q}=\{q\}$ [14, 19, 16, 15, 3] and $\mathcal{Q}=\{m,2m,\ldots,jm\}$ with $j,m>0$ [18, 2].
Recently, a bijection between the $k$-subsets of $\mathbb{N}_{n}$ with disallowed differences set $\mathcal{Q}$ of which the largest element is $q$ and the restricted-overlap tilings of an $(n+q)$-board (an $(n+q)\times 1$ board) with squares and $k$ combs was found. This enabled generating functions for $S_{n}$ and recursion relations for $S_{n,k}$ to be readily obtained for a variety of families of $\mathcal{Q}$ [1]. A $(w_{1},g_{1},w_{2},\ldots,g_{t-1},w_{t})$-comb is a tile composed of $t>0$ rectangular sub-tiles (called teeth) arranged in such a way that the $i$-th tooth, which has dimensions $w_{i}\times 1$, is separated from the $(i+1)$-th by a $g_{i}\times 1$ gap [1]. The length of a comb is the sum of the lengths of the teeth and gaps. Each $\mathcal{Q}$ corresponds to a length-$(q+1)$ comb $C$ (which we will refer to as a $\mathcal{Q}$-comb) defined as follows. We label the cells of $C$ from 0 to $q$. Cell 0 is always part of the first tooth. Cell $i$ (where $i=1,\ldots,q$) is part of a tooth iff $i\in\mathcal{Q}$. Note that we can always include 0 in $\mathcal{Q}$ (and we will do so occasionally for reasons of symmetry) since subset elements are distinct and so cannot differ by 0. This 0 element can then be regarded as corresponding to cell 0 of the comb. To ensure that there is only one way to specify the $w_{i}$ and $g_{i}$ of the $\mathcal{Q}$-comb, we insist that all teeth and gaps have positive integer length. A restricted-overlap tiling using squares and $\mathcal{Q}$-combs means that any comb can be placed so that its non-leftmost cells (i.e, cells 1 to $q$) overlap the non-leftmost cells of any other combs on the board and no other forms of overlap of tiles is allowed [1] (see Fig. 4 for examples). The following theorem is a consequence of the bijection.
Theorem 1 (Theorem 2 in [1]).
If $B^{\mathcal{Q}}_{n}$ is the number of restricted-overlap tilings of an $n$-board with $\mathcal{Q}$-combs and squares, and $B^{\mathcal{Q}}_{n,k}$ is the number of such tilings that contain $k$ combs, then $S^{\mathcal{Q}}_{n}=B^{\mathcal{Q}}_{n+q}$ and $S^{\mathcal{Q}}_{n,k}=B^{\mathcal{Q}}_{n+q,k}$.
Any tiling of an $n$-board can be expressed as a tiling using metatiles which are gapless groupings of tiles that cannot be split into smaller gapless groupings by removing a gapless grouping of tiles [9]. The simplest metatiles when tiling a long enough board with squares and combs are a single square (denoted by $S$) and a $(w_{1},g_{1},\ldots,w_{t})$-comb with all the gaps filled by squares. We call the latter metatile a filled comb and it is denoted by $CS^{g}$, where $g=\sum_{i}g_{i}$. If the comb has one tooth it is just a $w_{1}$-omino (and also a metatile by itself) and $S$ and $C$ are then the only metatiles. Otherwise, as a consequence of the restricted-overlap nature of the tiling, a $C$ can always be placed so that its leftmost cell occupies the next available empty cell of the yet-to-be-completed metatile. If more empty cells within the yet-to-be-completed metatile result (which, if a $C$ is placed starting in the final cell of the final gap of the first comb on the board, will be the case iff $2w_{t}<q$ – see Lemma 3.1/2 in [1]) this can continue indefinitely resulting in there being an infinite number of possible metatiles.
A systematic way to generate all possible metatiles and hence, in the simpler cases, obtain recursion relations for $B_{n}$ and $B_{n,k}$, is via a directed pseudograph (henceforth simply referred to as a digraph) in which each node has a label giving a bit-string representation of the remaining gaps (a 0 representing an empty cell, a 1 a filled one) starting at the next gap in the yet-to-be-completed metatile [10, 1] (see Figs. 1 and 2 for examples). The 0 node represents both the empty board and the completed metatile. Each arc represents the addition of a tile and has a label showing the type of tile and a subscript giving the increase in length of the incomplete metatile resulting from placing that tile, if that increase is not 0. The length increase from adding a $C$ is given by $q+1-d$, where $d$ is the number of binary digits in the node the arc starts from but with $d=0$ for the 0 node. A walk starting and ending at the 0 node but not visiting it in between represents a metatile and is denoted by the string of $C$ and $S$ corresponding to the arcs in the order visited. For example, in Fig. 1, the loop attached to the 0 node corresponds to the $S$ metatile and the cycle starting at the 0 node and then visiting the $10^{3}$ 1 node (an abbreviation for $0001$), the $10^{2}$ 1 node, and then the 01 node before returning to the 0 node corresponds to the $CS^{2}$ metatile. A cycle that does not include the 0 node is called an inner cycle (e.g., $SSC_{[4]}$ in Fig. 1 where including the length increment subscript quickly identifies which comb arc we are referring to). There are a finite number of possible metatiles iff the digraph lacks an inner cycle since a walk can traverse an inner cycle an arbitrary number of times before returning to the 0 node. If all the inner cycles have at least one node in common, that node is called a common node (e.g., the $01^{q-p}$ node is the common node in the Fig. 5 digraph; either the $10^{2}$ 1 or the 01 node can be chosen as the common node in the digraph in Fig. 2; Fig. 1 and Fig. 7 show digraphs that have inner cycles but no common node). For digraphs with a common node, as we review in §2, there are simple expressions for the recursion relations for numbers of tilings $B_{n}$ and $B_{n,k}$ in terms of lengths of various cycles in the digraph [10, 1]. If the digraph possesses inner cycles but no common node, as far as we are aware, there is only one other class of digraph (which has 3 inner cycles) where an analogous simple procedure has been found for obtaining the recursion relations [2].
A systematic search of the OEIS for sequences $(S^{\mathcal{Q}}_{n})_{n\geq 0}$ for all $\mathcal{Q}$ with $q\leq 4$ showed that there are various connections between some types of restricted combinations and some types of restricted permutations. Relations to compositions and bit strings were also found. Here, in §3, we give bijections between between two classes of strongly restricted permutations and combinations with $\mathcal{Q}=\{1,q\}$ and $\mathcal{Q}=\{m,2m,\ldots,jm\}$. As a consequence of the former bijection, we can prove a conjectured recursion relation for the permutations corresponding to the $\mathcal{Q}=\{1,4\}$ case. In §2, we derive recursion relations for $B_{n}$ and $B_{n,k}$ for tilings with metatiles generated by a particular class of digraph with 4 inner cycles but no common node. This enables one to also obtain recursion relations for $S^{\{1,5\}}_{n}$ and $S^{\{1,5\}}_{n,k}$. In §4, although it is not all that surprising given the tiling interpretation of restricted combinations, for completeness, we also show that $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into certain sets of parts for some classes of $\mathcal{Q}$. An elementary connection between all instances of $\mathcal{Q}$ and bit strings is pointed out in §5. This forms the basis for an algorithm for efficiently calculating $S_{n}$ and $S_{n,k}$ for any given $\mathcal{Q}$ as detailed via a heavily annotated C program listed in the Appendix. In §5 we also give a bijection between $k$-subsets for certain $\mathcal{Q}$ and equivalence classes for the occurrence of certain subwords within bit strings. The bijection enables us to show that a generating function for the number of equivalence classes is always straightforward to obtain.
2 Recursion relations for numbers of walks on 4-inner-cycle digraphs with two errant loops
We start by reviewing further terminology concerning digraphs. For a digraph possessing a common node, a walk from the 0 node to the common node followed by a walk from the common node back to the 0 node such that the common node is only visited once is called a common circuit (e.g., for either choice of common node, the digraph in Fig. 2 possesses a single common circuit, namely, $C_{[5]}S^{2}$). A cycle that includes the 0 node but not the common node is called an outer cycle (in Fig. 2, the $S$ loop attached to the 0 node is the only outer cycle).
A form of the following theorem was originally stated in [10]. A more compact version, re-derived in [1], is reproduced here. We will use this theorem in §3 to obtain the $S^{\{1,4\}}_{n}$ generating function.
Theorem 2 (Theorem 4 in [1]).
For a digraph possessing a common node, let $l_{\mathrm{o}i}$ be the length of the $i$-th outer cycle ($i=1,\ldots,N_{\rm o}$) and let $k_{\mathrm{o}i}$ be the number of combs it contains, let $L_{r}$ be the length of the $r$-th inner cycle ($r=1,\ldots,N$) and let $K_{r}$ be the number of combs it contains, and let $l_{\mathrm{c}i}$ be the length of the $i$-th common circuit ($i=1,\ldots,N_{\rm c}$) and let $k_{\mathrm{c}i}$ be the number of combs it contains. Then for all integers $n$ and $k$,
$B_{n}$ $=\delta_{n,0}+\sum_{r=1}^{N}(B_{n-L_{r}}-\delta_{n,L_{r}})+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i}}-\sum_{r=1}^{N}B_{n-l_{\mathrm{o}i}-L_{r}}\biggr{)}+\sum_{i=1}^{N_{\rm c}}B_{n-l_{\mathrm{c}i}},$
$B_{n,k}$ $=\delta_{n,0}\delta_{k,0}+\sum_{r=1}^{N}(B_{n-L_{r},k-K_{r}}-\delta_{n,L_{r}}\delta_{k,K_{r}})+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i},k-k_{\mathrm{o}i}}-\sum_{r=1}^{N}B_{n-l_{\mathrm{o}i}-L_{r},k-k_{\mathrm{o}i}-K_{r}}\biggr{)}$
$\qquad\mbox{}+\sum_{i=1}^{N_{\rm c}}B_{n-l_{\mathrm{c}i},k-k_{\mathrm{c}i}},$
where $B_{n<0}=B_{n,k<0}=B_{n<k,k}=0$.
The following definitions are slightly modified versions of those given in earlier work on a class of digraph with inner cycles but no common node in order to incorporate the increased complexity of the type of digraphs we are about to consider [2]. For a digraph lacking a common node, we refer to any inner cycle that can be represented as a single arc linking a node to itself as an errant loop if the digraph would only have a common node $\mathcal{P}$ if all the errant loop arcs were removed. The node $\mathcal{P}$ is then referred to as a pseudo-common node. A node with an errant loop is called an errant loop node. Evidently, a particular errant loop node $\mathcal{E}$ cannot also be a pseudo-common node; if that were the case, removing all the other errant loop arcs except the errant loop at $\mathcal{E}$ would result in a digraph with a common node implying that the loop at $\mathcal{E}$ was never an errant loop in the first place. For a digraph with at least one errant loop, a common circuit is now defined as two concatenated simple paths from the 0 node to $\mathcal{P}$ and from $\mathcal{P}$ to the 0 node. We also need to modify the definition of an outer cycle: it is now a cycle that includes the 0 node but not $\mathcal{P}$.
The proof of Theorem 6 requires the following three results, the first of which is the $N=2$ case of Lemma 35 in [2].
Lemma 3.
For positive integers $j_{1},j_{2},j_{4}$ we have
$\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{2}+j_{4}-1}{j_{4}}\!=\!\binom{j_{1}+j_{2}-1}{j_{1}-1}\Biggl{(}\!\binom{j_{2}+j_{4}-1}{j_{4}}-\binom{j_{2}+j_{4}-2}{j_{4}-1}\!\Biggr{)}\\ +\binom{j_{1}+j_{2}-1}{j_{1}}\binom{j_{2}+j_{4}-2}{j_{4}}+\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{2}+j_{4}-2}{j_{4}-1}.$ (1)
Corollary 4.
For positive integers $j_{1},j_{2},j_{3}$ we have
$\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-1}{j_{3}}\!=\!\binom{j_{1}+j_{2}-1}{j_{1}}\Biggl{(}\!\binom{j_{1}+j_{3}-1}{j_{3}}-\binom{j_{1}+j_{3}-2}{j_{3}-1}\!\Biggr{)}\\ +\binom{j_{1}+j_{2}-1}{j_{1}-1}\binom{j_{1}+j_{3}-2}{j_{3}}+\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-2}{j_{3}-1}.$ (2)
Proof.
Change $j_{1},j_{2},j_{4}$ in Lemma 3 to $j_{2},j_{1},j_{3}$, respectively, and then re-express $\tbinom{j_{1}+j_{2}-1}{j_{2}-1}$ as $\tbinom{j_{1}+j_{2}-1}{j_{1}}$ and $\tbinom{j_{1}+j_{2}-1}{j_{2}}$ as $\tbinom{j_{1}+j_{2}-1}{j_{1}-1}$. ∎
Lemma 5.
For positive integers $j_{i}$ ($i=1,2,3,4$) we have
$\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-1}{j_{3}}\binom{j_{2}+j_{4}-1}{j_{4}}\\ =\binom{j_{1}+j_{2}-1}{j_{1}-1}\binom{j_{1}+j_{3}-2}{j_{3}}\binom{j_{2}+j_{4}-1}{j_{4}}+\binom{j_{1}+j_{2}-1}{j_{1}}\binom{j_{1}+j_{3}-1}{j_{3}}\binom{j_{2}+j_{4}-2}{j_{4}}\\ +\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-2}{j_{3}-1}\binom{j_{2}+j_{4}-1}{j_{4}}+\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-1}{j_{3}}\binom{j_{2}+j_{4}-2}{j_{4}-1}\\ -\binom{j_{1}+j_{2}-1}{j_{1}-1}\binom{j_{1}+j_{3}-2}{j_{3}}\binom{j_{2}+j_{4}-2}{j_{4}-1}-\binom{j_{1}+j_{2}-1}{j_{1}}\binom{j_{1}+j_{3}-2}{j_{3}-1}\binom{j_{2}+j_{4}-2}{j_{4}}\\ -\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-2}{j_{3}-1}\binom{j_{2}+j_{4}-2}{j_{4}-1}.$ (3)
Proof.
The right-hand side of (3), after combining the first term with the fifth, the second term with the sixth, and the fourth term with the seventh, can be re-expressed as
$\Biggl{(}\binom{j_{1}+j_{2}-1}{j_{1}-1}+\binom{j_{1}+j_{2}-1}{j_{1}}\Biggr{)}\binom{j_{1}+j_{3}-2}{j_{3}}\binom{j_{2}+j_{4}-2}{j_{4}}$
$\qquad+\binom{j_{1}+j_{2}}{j_{1}}\Biggl{(}\binom{j_{1}+j_{3}-2}{j_{3}-1}\binom{j_{2}+j_{4}-1}{j_{4}}+\binom{j_{1}+j_{3}-2}{j_{3}}\binom{j_{2}+j_{4}-2}{j_{4}-1}\Biggr{)}$
$=\binom{j_{1}+j_{2}}{j_{1}}\Biggl{(}\binom{j_{1}+j_{3}-2}{j_{3}-1}\binom{j_{2}+j_{4}-1}{j_{4}}+\binom{j_{1}+j_{3}-2}{j_{3}}\binom{j_{2}+j_{4}-1}{j_{4}}\Biggr{)},$
which can be seen to reduce to the left-hand side of (3). ∎
Theorem 6.
If a digraph has outer cycles of lengths $l_{\mathrm{o}i}$ that contain $k_{\mathrm{o}i}$ combs ($i=1\ldots,N_{\rm o}$), and errant loops of lengths $L_{3}$ and $L_{4}$ containing $K_{3}$ and $K_{4}$ combs, respectively, attached to inner cycles of lengths $L_{1}$ and $L_{2}$ containing $K_{1}$ and $K_{2}$ combs, respectively, and the outgoing part of the $i$-th of these inner cycles is part of the $i$-th common circuit which is of length $l_{\mathrm{c}i}$ and contains $k_{\mathrm{c}i}$ combs ($i=1,2$), then for all integers $n$ and $k$ we have
$B_{n}$ $=\delta_{n,0}+\sum_{r=1}^{4}(B_{n-L_{r}}-\delta_{n,L_{r}})$
$\qquad+\delta_{n,L_{1}+L_{4}}-B_{n-L_{1}-L_{4}}+\delta_{n,L_{2}+L_{3}}-B_{n-L_{2}-L_{3}}+\delta_{n,L_{3}+L_{4}}-B_{n-L_{3}-L_{4}}$
$\qquad+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i}}+B_{n-l_{\mathrm{o}i}-L_{1}-L_{4}}+B_{n-l_{\mathrm{o}i}-L_{2}-L_{3}}+B_{n-l_{\mathrm{o}i}-L_{3}-L_{4}}-\sum_{r=1}^{4}B_{n-l_{\mathrm{o}i}-L_{r}}\biggr{)}$
$\qquad+\sum_{i=1}^{2}(B_{n-l_{\mathrm{c}i}}-B_{n-l_{\mathrm{c}i}-L_{5-i}}),$ (4)
$B_{n,k}$ $=\delta_{n,0}\delta_{k,0}+\sum_{r=1}^{4}(B_{n-L_{r},k-K_{r}}-\delta_{n,L_{r}}\delta_{k,K_{r}})+\delta_{n,L_{1}+L_{4}}\delta_{k,K_{1}+K_{4}}-B_{n-L_{1}-L_{4},k-K_{1}-K_{4}}$
$\qquad+\delta_{n,L_{2}+L_{3}}\delta_{k,K_{2}+K_{3}}-B_{n-L_{2}-L_{3},k-K_{2}-K_{3}}+\delta_{n,L_{3}+L_{4}}\delta_{k,K_{3}+K_{4}}-B_{n-L_{3}-L_{4},k-K_{3}-K_{4}}$
$\qquad+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i},k-k_{\mathrm{o}i}}-\sum_{r=1}^{4}B_{n-l_{\mathrm{o}i}-L_{r},k-k_{\mathrm{o}i}-K_{r}}+B_{n-l_{\mathrm{o}i}-L_{1}-L_{4},k-k_{\mathrm{o}i}-K_{1}-K_{4}}$
$\qquad\qquad+B_{n-l_{\mathrm{o}i}-L_{2}-L_{3},k-k_{\mathrm{o}i}-K_{2}-K_{3}}+B_{n-l_{\mathrm{o}i}-L_{3}-L_{4},k-k_{\mathrm{o}i}-K_{3}-K_{4}}\biggr{)}$
$\qquad+\sum_{i=1}^{2}(B_{n-l_{\mathrm{c}i},k-k_{\mathrm{c}i}}-B_{n-l_{\mathrm{c}i}-L_{5-i},k-k_{\mathrm{c}i}-K_{5-i}}),$ (5)
where $B_{n<0}=B_{n,k<0}=B_{n<k,k}=0$.
Figure 1: Digraph for restricted-overlap tiling with squares ($S$) and $(2,3,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,5\}$ case.
Proof.
We begin by proving (4) by conditioning on the final metatile on the board [8]. If it is of length $l$ then there are $B_{n-l}$ ways to tile the remainder of the board. If $n=l$ then there is just one way to tile the board with that metatile which necessitates choosing $B_{0}$ to be 1 which is the reason for including the $\delta_{n,0}$ term. Considering all possible metatiles leads to
$B_{n}=\delta_{n,0}+\!\sum_{i=1}^{N_{\rm o}}B_{n-l_{\mathrm{o}i}}+\!\sum_{i=1}^{2}\sum_{e=0}^{\infty}\Biggl{(}\beta_{0}+\!\!\sum_{\begin{subarray}{c}j_{1}\geq 1,\\ j_{3}\geq 0\end{subarray}}\!\binom{j_{1}+j_{3}-1}{j_{3}}\beta_{j_{1}L_{1}+j_{3}L_{3}}+\!\!\sum_{\begin{subarray}{c}j_{2}\geq 1,\\ j_{4}\geq 0\end{subarray}}\!\binom{j_{2}+j_{4}-1}{j_{4}}\beta_{j_{2}L_{2}+j_{4}L_{4}}\\ +\sum_{\begin{subarray}{c}j_{1},j_{2}\geq 1,\\ j_{3},j_{4}\geq 0\end{subarray}}\!\!\!\binom{j_{1}+j_{2}}{j_{1}}\binom{j_{1}+j_{3}-1}{j_{3}}\binom{j_{2}+j_{4}-1}{j_{4}}\beta_{j_{1}L_{1}+j_{2}L_{2}+j_{3}L_{3}+j_{4}L_{4}}\Biggr{)},$ (6)
where $\beta_{a}=B_{n-l_{\mathrm{c}i}-eL_{i+2}-a}$. We now explain the origin of the various terms inside the sums and in each case give an example of a corresponding family of metatiles from the digraph in Fig. 1, taking $10^{3}$ 1 to be the pseudo-common node $\mathcal{P}$. The $\beta_{0}$ term corresponds to traversing the $i$-th common circuit up to its errant node, executing the errant loop $e$ times, and then completing the circuit (e.g., $C_{[6]}SC_{[3]}^{e}S^{2}$). The remaining terms also start with the part of the $i$-th common circuit up to $\mathcal{P}$ and end with the remainder of the $i$-th common circuit from $\mathcal{P}$ with $e$ repetitions of the corresponding errant loop. The first of the sums corresponds to the first (length $L_{1}$) inner cycle getting executed $j_{1}\geq 1$ times and its errant loop (of length $L_{3}$) also getting executed $j_{3}\geq 0$ times in any order except that the first inner cycle must be entered before its corresponding errant loop, which accounts for the $\tbinom{j_{1}+j_{3}-1}{j_{3}}$ (rather than $\tbinom{j_{1}+j_{3}}{j_{3}}$) factor (e.g., $C_{[6]}S^{2}C_{[4]}SC_{[3]}^{2}SC_{[4]}C_{[2]}^{e+1}S^{2}$). The second sum is analogous to the first but this time applies to the second (length $L_{2}$) inner cycle and its corresponding (length $L_{4}$) errant loop (e.g., $C_{[6]}C_{[2]}^{3}SC_{[4]}C_{[2]}^{e+1}S^{2}$). The final sum corresponds to the first two inner cycles getting executed $j_{1}$ and $j_{2}$ times, respectively, in any order (which gives the $\binom{j_{1}+j_{2}}{j_{1}}$ factor), and while executing those cycles, the errant loops can be visited an arbitrary number of times (e.g., $C_{[6]}C_{[2]}SC_{[4]}S^{2}C_{[4]}SC_{[4]}SC_{[3]}^{3}SC_{[4]}SC_{[3]}^{e}S^{2}$).
Note that in the above explanations we have assumed that $\mathcal{P}$ lies before the errant node in each common circuit. The theorem still applies if $\mathcal{P}$ occurs after the errant node in each common circuit (which would be case had we selected $\mathcal{P}$ to be the $01$ node in Fig. 1).
To keep the expressions in the remainder of the proof to a manageable size, we introduce the following notation: $\lambda_{a}=j_{a}L_{a}$, $\lambda_{ab}=j_{a}L_{a}+j_{b}L_{b}$, $\lambda_{abc}=j_{a}L_{a}+j_{b}L_{b}+j_{c}L_{c}$, $\lambda=\sum_{r=1}^{4}j_{r}L_{r}$,
$\sum_{abcd}\text{~{}means}\!\!\!\!\sum_{\begin{subarray}{c}j_{1}\geq a,\,j_{2}\geq b\\ j_{3}\geq c,\,j_{4}\geq d\end{subarray}},\qquad\begin{vmatrix}a&b&c\\ d&e&f\end{vmatrix}=\binom{j_{1}+j_{2}-a}{j_{1}-d}\binom{j_{1}+j_{3}-b}{j_{3}-e}\binom{j_{2}+j_{4}-c}{j_{4}-f},$
and if a binomial coefficient or sum over an index is not present, then a : is placed in the corresponding position. For example, after rearranging and then using this notation, (3), (1), and (2) can be re-expressed, respectively, as
$\begin{vmatrix}0&1&1\\ 0&0&0\end{vmatrix}+\begin{vmatrix}1&2&2\\ 1&0&1\end{vmatrix}+\begin{vmatrix}1&2&2\\ 0&1&0\end{vmatrix}+\begin{vmatrix}0&2&2\\ 0&1&1\end{vmatrix}=\begin{vmatrix}1&2&1\\ 1&0&0\end{vmatrix}+\begin{vmatrix}1&1&2\\ 0&0&0\end{vmatrix}+\begin{vmatrix}0&2&1\\ 0&1&0\end{vmatrix}+\begin{vmatrix}0&1&2\\ 0&0&1\end{vmatrix},$ (7)
$\begin{vmatrix}0&:&1\\ 0&:&0\end{vmatrix}+\begin{vmatrix}1&:&2\\ 1&:&1\end{vmatrix}=\begin{vmatrix}1&:&1\\ 1&:&0\end{vmatrix}+\begin{vmatrix}1&:&2\\ 0&:&0\end{vmatrix}+\begin{vmatrix}0&:&2\\ 0&:&1\end{vmatrix},$ (8)
$\begin{vmatrix}0&1&:\\ 0&0&:\end{vmatrix}+\begin{vmatrix}1&2&:\\ 0&1&:\end{vmatrix}=\begin{vmatrix}1&1&:\\ 0&0&:\end{vmatrix}+\begin{vmatrix}1&2&:\\ 1&0&:\end{vmatrix}+\begin{vmatrix}0&2&:\\ 0&1&:\end{vmatrix},$ (9)
and the three sums enclosed by the largest parentheses in (6) become
$\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}+\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}}+\sum_{1100}\begin{vmatrix}0&1&1\\ 0&0&0\end{vmatrix}\beta_{\lambda}.$
Representing (6) by $E(n)$, we then write down
$E(n)-\sum_{r=1}^{4}E(n-L_{r})+E(n-L_{1}-L_{4})+E(n-L_{2}-L_{3})+E(n-L_{3}-L_{4})$
which, after re-indexing some of the sums so that the subscripts of $\beta$ remain the same as in $E(n)$, yields
$B_{n}-\sum_{r=1}^{4}B_{n-L_{r}}+B_{n-L_{1}-L_{4}}+B_{n-L_{2}-L_{3}}+B_{n-L_{3}-L_{4}}=\delta_{n,0}-\sum_{r=1}^{4}\delta_{n,L_{r}}+\delta_{n,L_{1}+L_{4}}+\delta_{n,L_{2}+L_{3}}$
$\quad+\delta_{n,L_{3}+L_{4}}+\sum_{i=1}^{N_{\rm o}}\biggl{(}B_{n-l_{\mathrm{o}i}}-\sum_{r=1}^{4}B_{n-l_{\mathrm{o}i}-L_{r}}+B_{n-l_{\mathrm{o}i}-L_{1}-L_{4}}+B_{n-l_{\mathrm{o}i}-L_{2}-L_{3}}+B_{n-l_{\mathrm{o}i}-L_{3}-L_{4}}\biggr{)}$
$\quad+\sum_{i=1}^{2}\sum_{e=0}^{\infty}\Biggl{(}\beta_{0}-\beta_{L_{1}}-\beta_{L_{2}}-\beta_{L_{3}}-\beta_{L_{4}}+\beta_{L_{1}+L_{4}}+\beta_{L_{2}+L_{3}}+\beta_{L_{3}+L_{4}}$
$\qquad\mbox{}+\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}-\sum_{2:0:}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}-\sum_{1:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}}-\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}-\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}$
$\qquad\mbox{}+\sum_{2:0:}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}+\sum_{1:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}+\sum_{1:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}$
$\qquad\mbox{}+\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}}-\sum_{:2:0}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}}-\sum_{:1:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}}-\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{1}}-\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{3}}$
$\qquad\mbox{}+\sum_{:2:0}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{3}}+\sum_{:1:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{1}}+\sum_{:1:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{3}}$
$\qquad\mbox{}+\sum_{1100}\begin{vmatrix}0&1&1\\ 0&0&0\end{vmatrix}\beta_{\lambda}-\sum_{2100}\begin{vmatrix}1&2&1\\ 1&0&0\end{vmatrix}\beta_{\lambda}-\sum_{1200}\begin{vmatrix}1&1&2\\ 0&0&0\end{vmatrix}\beta_{\lambda}-\sum_{1110}\begin{vmatrix}0&2&1\\ 0&1&0\end{vmatrix}\beta_{\lambda}-\sum_{1101}\begin{vmatrix}0&1&2\\ 0&0&1\end{vmatrix}\beta_{\lambda}$
$\qquad\mbox{}+\sum_{2101}\begin{vmatrix}1&2&2\\ 1&0&1\end{vmatrix}\beta_{\lambda}+\sum_{1210}\begin{vmatrix}1&2&2\\ 0&1&0\end{vmatrix}\beta_{\lambda}+\sum_{1111}\begin{vmatrix}0&2&2\\ 0&1&1\end{vmatrix}\beta_{\lambda}\Biggr{)}.$ (10)
Rearranging the first two lines of (10) gives the first three lines of (4). The sum over $e$ of $\beta_{0}-\beta_{L_{3}}-\beta_{L_{4}}+\beta_{L_{3}+L_{4}}$ for $i=1,2$ can be written as
$\sum_{e=0}^{\infty}B_{n-l_{\mathrm{c}i}-eL_{i+2}}-B_{n-l_{\mathrm{c}i}-(e+1)L_{i+2}}-(B_{n-l_{\mathrm{c}i}-eL_{i+2}-L_{5-i}}-B_{n-l_{\mathrm{c}i}-(e+1)L_{i+2}-L_{5-i}}),$
which, on expanding the sum, leaves $B_{n-l_{\mathrm{c}i}}-B_{n-l_{\mathrm{c}i}-L_{5-i}}$ and hence the last two terms in (4) are accounted for. We now show that all the remaining terms in (10) cancel out by first re-expressing each of the 24 sums inside the double sum.
$\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}=\sum_{2:1:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}+\sum_{2:::}\beta_{\lambda_{1}}+\sum_{::1:}\beta_{\lambda_{3}+L_{1}}+\beta_{L_{1}},$ (11a)
$\sum_{2:0:}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}=\sum_{2:1:}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}}+\sum_{2:::}\beta_{\lambda_{1}},$ (11b)
$\sum_{1:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}}=\sum_{2:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}}+\sum_{::1:}\beta_{\lambda_{3}+L_{1}},$ (11c)
$\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}=\sum_{2:1:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}+\sum_{2:::}\beta_{\lambda_{1}+L_{2}}+\sum_{::1:}\beta_{\lambda_{3}+L_{1}+L_{2}}+\beta_{L_{1}+L_{2}},$ (11d)
$\sum_{1:0:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}=\sum_{2:1:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}+\sum_{2:::}\beta_{\lambda_{1}+L_{4}}+\sum_{::1:}\beta_{\lambda_{3}+L_{1}+L_{4}}+\beta_{L_{1}+L_{4}},$ (11e)
$\sum_{2:0:}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}=\sum_{2:1:}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}+\sum_{2:::}\beta_{\lambda_{1}+L_{4}},$ (11f)
$\sum_{1:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}=\sum_{2:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}+\sum_{::1:}\beta_{\lambda_{3}+L_{1}+L_{2}},$ (11g)
$\sum_{1:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}=\sum_{2:1:}\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{4}}+\sum_{::1:}\beta_{\lambda_{3}+L_{1}+L_{4}},$ (11h)
$\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{13}}=\sum_{:2:1}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}}+\sum_{:2::}\beta_{\lambda_{2}}+\sum_{:::1}\beta_{\lambda_{4}+L_{2}}+\beta_{L_{2}},$ (11i)
$\sum_{:2:0}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}}=\sum_{:2:1}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}}+\sum_{:2::}\beta_{\lambda_{2}},$ (11j)
$\sum_{:1:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}}=\sum_{:2:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}}+\sum_{:::1}\beta_{\lambda_{4}+L_{2}},$ (11k)
$\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{1}}=\sum_{:2:1}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{1}}+\sum_{:2::}\beta_{\lambda_{2}+L_{1}}+\sum_{:::1}\beta_{\lambda_{4}+L_{1}+L_{2}}+\beta_{L_{1}+L_{2}},$ (11l)
$\sum_{:1:0}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{3}}=\sum_{:2:1}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{3}}+\sum_{:2::}\beta_{\lambda_{2}+L_{3}}+\sum_{:::1}\beta_{\lambda_{4}+L_{2}+L_{3}}+\beta_{L_{2}+L_{3}},$ (11m)
$\sum_{:2:0}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{3}}=\sum_{:2:1}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{3}}+\sum_{:2::}\beta_{\lambda_{2}+L_{3}},$ (11n)
$\sum_{:1:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{1}}=\sum_{:2:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{1}}+\sum_{:::1}\beta_{\lambda_{4}+L_{1}+L_{2}},$ (11o)
$\sum_{:1:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{3}}=\sum_{:2:1}\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{3}}+\sum_{:::1}\beta_{\lambda_{4}+L_{2}+L_{3}},$ (11p)
$\sum_{1100}\begin{vmatrix}0&1&1\\ 0&0&0\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}0&1&1\\ 0&0&0\end{vmatrix}\beta_{\lambda}+\sum_{:211}(j_{2}+1)\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{234}+L_{1}}+\sum_{2:11}(j_{1}+1)\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{134}+L_{2}}$
$\quad+\sum_{22:1}\begin{vmatrix}0&:&1\\ 0&:&0\end{vmatrix}\beta_{\lambda_{124}}+\sum_{221:}\begin{vmatrix}0&1&:\\ 0&0&:\end{vmatrix}\beta_{\lambda_{123}}+2\sum_{::11}\beta_{\lambda_{34}+L_{1}+L_{2}}+\sum_{22::}\begin{vmatrix}0&:&:\\ 0&:&:\end{vmatrix}\beta_{\lambda_{12}}$
$\quad+\sum_{2:1:}(j_{1}+1)\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}+\sum_{:2:1}(j_{2}+1)\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{1}}$
$\quad+\sum_{2::1}(j_{1}+1)\beta_{\lambda_{14}+L_{2}}+\sum_{:21:}(j_{2}+1)\beta_{\lambda_{23}+L_{1}}+\sum_{2:::}(j_{1}+1)\beta_{\lambda_{1}+L_{2}}+\sum_{:2::}(j_{2}+1)\beta_{\lambda_{2}+L_{1}}$
$\quad+2\sum_{::1:}\beta_{\lambda_{3}+L_{1}+L_{2}}+2\sum_{:::1}\beta_{\lambda_{4}+L_{1}+L_{2}}+2\beta_{L_{1}+L_{2}},$ (11q)
$\sum_{2100}\begin{vmatrix}1&2&1\\ 1&0&0\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}1&2&1\\ 1&0&0\end{vmatrix}\beta_{\lambda}+\sum_{2:11}j_{1}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{134}+L_{2}}+\sum_{22:1}\begin{vmatrix}1&:&1\\ 1&:&0\end{vmatrix}\beta_{\lambda_{124}}+\sum_{221:}\begin{vmatrix}1&2&:\\ 1&0&:\end{vmatrix}\beta_{\lambda_{123}}$
$\quad+\sum_{22::}\begin{vmatrix}1&:&:\\ 1&:&:\end{vmatrix}\beta_{\lambda_{12}}+\sum_{2:1:}j_{1}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}+\sum_{2::1}j_{1}\beta_{\lambda_{14}+L_{2}}+\sum_{2:::}j_{1}\beta_{\lambda_{1}+L_{2}},$ (11r)
$\sum_{1200}\begin{vmatrix}1&1&2\\ 0&0&0\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}1&1&2\\ 0&0&0\end{vmatrix}\beta_{\lambda}+\sum_{:211}j_{2}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{234}+L_{1}}+\sum_{22:1}\begin{vmatrix}1&:&2\\ 0&:&0\end{vmatrix}\beta_{\lambda_{124}}+\sum_{221:}\begin{vmatrix}1&1&:\\ 0&0&:\end{vmatrix}\beta_{\lambda_{123}}$
$\quad+\sum_{22::}\begin{vmatrix}1&:&:\\ 0&:&:\end{vmatrix}\beta_{\lambda_{12}}+\sum_{:2:1}j_{2}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24}+L_{1}}+\sum_{:21:}j_{2}\beta_{\lambda_{23}+L_{1}}+\sum_{:2::}j_{2}\beta_{\lambda_{2}+L_{1}},$ (11s)
$\sum_{1110}\begin{vmatrix}0&2&1\\ 0&1&0\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}0&2&1\\ 0&1&0\end{vmatrix}\beta_{\lambda}$
$\quad+\sum_{221:}\begin{vmatrix}0&2&:\\ 0&1&:\end{vmatrix}\beta_{\lambda_{123}}+\sum_{2:11}(j_{1}+1)\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{134}+L_{2}}+\sum_{:211}(j_{2}+1)\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{234}+L_{1}}$
$\quad+\sum_{2:1:}(j_{1}+1)\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{13}+L_{2}}+\sum_{:21:}(j_{2}+1)\beta_{\lambda_{23}+L_{1}}+2\sum_{::11}\beta_{\lambda_{34}+L_{1}+L_{2}}+2\sum_{::1:}\beta_{\lambda_{3}+L_{1}+L_{2}},$ (11t)
$\sum_{1101}\begin{vmatrix}0&1&2\\ 0&0&1\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}0&1&2\\ 0&0&1\end{vmatrix}\beta_{\lambda}$
$\quad+\sum_{22:1}\begin{vmatrix}0&:&2\\ 0&:&1\end{vmatrix}\beta_{\lambda_{124}}+\sum_{2:11}(j_{1}+1)\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{134}+L_{2}}+\sum_{:211}(j_{2}+1)\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{234}+L_{1}}$
$\quad+\sum_{:2:1}(j_{2}+1)\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{24}+L_{1}}+\sum_{2::1}(j_{1}+1)\beta_{\lambda_{14}+L_{2}}+2\sum_{::11}\beta_{\lambda_{34}+L_{1}+L_{2}}+2\sum_{:::1}\beta_{\lambda_{4}+L_{1}+L_{2}},$ (11u)
$\sum_{2101}\begin{vmatrix}1&2&2\\ 1&0&1\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}1&2&2\\ 1&0&1\end{vmatrix}\beta_{\lambda}+\sum_{22:1}\begin{vmatrix}1&:&2\\ 1&:&1\end{vmatrix}\beta_{\lambda_{124}}+\sum_{2:11}j_{1}\begin{vmatrix}:&2&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{134}+L_{2}}+\sum_{2::1}j_{1}\beta_{\lambda_{14}+L_{2}},$ (11v)
$\sum_{1210}\begin{vmatrix}1&2&2\\ 0&1&0\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}1&2&2\\ 0&1&0\end{vmatrix}\beta_{\lambda}+\sum_{221:}\begin{vmatrix}1&2&:\\ 0&1&:\end{vmatrix}\beta_{\lambda_{123}}+\sum_{:211}j_{2}\begin{vmatrix}:&:&2\\ :&:&0\end{vmatrix}\beta_{\lambda_{234}+L_{1}}+\sum_{:21:}j_{2}\beta_{\lambda_{23}+L_{1}},$ (11w)
$\sum_{1111}\begin{vmatrix}0&2&2\\ 0&1&1\end{vmatrix}\beta_{\lambda}=\sum_{2211}\begin{vmatrix}0&2&2\\ 0&1&1\end{vmatrix}\beta_{\lambda}+\sum_{2:11}(j_{1}+1)\begin{vmatrix}:&2&:\\ :&1&:\end{vmatrix}\beta_{\lambda_{134}+L_{2}}+\sum_{:211}(j_{2}+1)\begin{vmatrix}:&:&2\\ :&:&1\end{vmatrix}\beta_{\lambda_{234}+L_{1}}$
$\quad+2\sum_{::11}\beta_{\lambda_{34}+L_{1}+L_{2}}.$ (11x)
[…]
The proof of (5) proceeds in an analogous way. Conditioning on the final metatile gives
$B_{n,k}=\delta_{n,0}\delta_{k,0}+\!\sum_{i=1}^{N_{\rm o}}B_{n-l_{\mathrm{o}i},k-k_{\mathrm{o}i}}+\!\sum_{i=1}^{2}\sum_{e=0}^{\infty}\Biggl{(}\beta_{0,0}+\!\!\sum_{1:1:}\begin{vmatrix}:&1&:\\ :&0&:\end{vmatrix}\beta_{\lambda_{13},\kappa_{13}}+\!\!\sum_{:1:1}\begin{vmatrix}:&:&1\\ :&:&0\end{vmatrix}\beta_{\lambda_{24},\kappa_{24}}\\ +\sum_{1100}\begin{vmatrix}0&1&1\\ 0&0&0\end{vmatrix}\beta_{\lambda,\kappa}\Biggr{)},$ (12)
with $B_{n,k>n}=B_{n,k<0}=0$ and where $\beta_{a,b}=B_{n-l_{\mathrm{c}i}-eL_{i+2}-a,k-k_{\mathrm{c}i}-eK_{i+2}-b}$, $\kappa_{ab}=j_{a}K_{a}+j_{b}K_{b}$, and $\kappa=\sum_{r=1}^{4}j_{r}K_{r}$. Denoting (12) by $E(n,k)$, writing down
$E(n,k)-\sum_{r=1}^{4}E(n-L_{r},k-K_{r})+E(n-L_{1}-L_{4},k-K_{1}-K_{4})+E(n-L_{2}-L_{3},k-K_{2}-K_{3})\\ +E(n-L_{3}-L_{4},k-K_{3}-K_{4}),$
and then proceeding in the same way as for the proof of (4) gives the required result. ∎
3 Bijections between strongly restricted permutations and restricted combinations
A strongly restricted permutation $\pi$ of the set $\mathbb{N}_{n}$ is a permutation for which the number of permissible values of $\pi(i)-i$ for each $i\in\mathbb{N}_{n}$ is less than a finite number independent of $n$ [17]. Here we give two bijections between restricted combinations and types of strongly restricted permutations where the permutations satisfy $\pi(i)-i\in\mathcal{D}$. Our first bijection concerns a strongly restricted permutation where $\mathcal{D}=\{-1,0,-1\}$ and any $m$ consecutive $\pi(i)$ cannot differ from one another by more than $m$.
Theorem 7.
There is a bijection between the $k$-subsets of $\mathbb{N}_{n}$ such that no two elements of the subset differ by an element of $\mathcal{Q}=\{1,m\}$, where $m=2,3,\ldots$, and the permutations $\pi$ of $\mathbb{N}_{n+1}$ that have $k$ pairs of digits exchanged such that (a) the pairs do not overlap and (b) for all $i=1,\ldots,n+2-m$ the permutation satisfies $\max_{j=0,\ldots,m-1}\pi(i+j)-\min_{j=0,\ldots,m-1}\pi(i+j)\leq m$.
Proof.
The permutation $\pi$ of $\mathbb{N}_{n+1}$ corresponding to a subset $\mathcal{S}$ of $\mathbb{N}_{n}$ is formed as follows. If $i\in\mathcal{S}$ (where $i=1,\ldots,n$) then $\pi(i)=i+1$ and $\pi(i+1)=i$ (i.e., a pair of exchanged digits) whereas if neither $i$ nor $i-1$ are in $\mathcal{S}$ (for $i=2,\ldots,n+1$) then $\pi(i)=i$. As $\mathcal{Q}$ contains 1, the pairs of exchanged digits in the permutation cannot overlap. Given that condition (a) is satisfied, the only way for (b) to be violated would be if $\pi(i+1)=i$ and $\pi(i+m)=i+m+1$ but this could only occur if both $i$ and $i+m$ were in $\mathcal{S}$ which is impossible since $m\in\mathcal{Q}$. The process of forming a permutation from a subset is clearly reversible and so the bijection is established. ∎
Theorem 7 has connections to the following sequences in the OEIS [20]: A000930 ($S^{\{1,2\}}_{n}$) and A102547 ($S^{\{1,2\}}_{n,k}$), its associated triangle; A130137 ($S^{\{1,3\}}_{n}$), which we mention again in the section on bit strings; A263710 ($S^{\{1,4\}}_{n}$), which was the sequence that led to the discovery of the bijection and we will now prove the conjectured generating function for that entry; and A374737 ($S^{\{1,5\}}_{n}$), which was recently added by the author.
The case $\mathcal{Q}=\{1,4\}$ falls into the class covered by the final theorem in [1], which gives recursion relations for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$. However, in order to avoid introducing the notation that theorem requires at this stage, we instead obtain the generating function using Theorem 2 which is also straightforward and illustrative to do.
Figure 2: Digraph for restricted-overlap tiling with squares ($S$) and $(2,2,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,4\}$ case.
Proposition 8.
The generating function $G(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,4\}$ is given by
$G(x)=\frac{1+x+x^{2}+x^{3}+2x^{4}+x^{6}}{1-x-x^{3}+x^{4}-2x^{5}+x^{6}-x^{7}}.$ (13)
Proof.
The $\mathcal{Q}$-comb is a $(2,2,1)$-comb. The metatile-generating digraph for tiling with squares and such combs (Fig. 2) has two inner cycles ($SC_{[3]}$ and $C_{[2]}C_{[3]}$) which have lengths $L_{1}=3$ and $L_{2}=5$. We choose $10^{2}$ 1 as the common node. The common circuits ($C_{[5]}S^{2}$ and $C_{[5]}C_{[2]}S$) have lengths $l_{\rm c1}=5$ and $l_{\rm c2}=7$. There is a single outer cycle ($S$) which has length $l_{\rm o1}=1$. Using these in Theorem 2 yields
$B_{n}=\delta_{n,0}-\delta_{n,3}-\delta_{n,5}+B_{n-1}+B_{n-3}-B_{n-4}+2B_{n-5}-B_{n-6}+B_{n-7}.$ (14)
The generating function for $S_{n}$ is therefore
$\frac{1}{x^{4}}\biggl{(}\frac{1-x^{3}-x^{5}}{1-x-x^{3}+x^{4}-2x^{5}+x^{6}-x^{7}}-\frac{1-x^{4}}{1-x}\biggr{)},$
which reduces to (13). ∎
As our first application of Theorem 6, we also obtain the generating function for the next sequence in the series, i.e., that corresponding to $\mathcal{Q}=\{1,5\}$ (A374737).
Proposition 9.
The generating function $G(x)$ for the number of restricted combinations of $\mathbb{N}_{n}$ with disallowed differences $\{1,5\}$ is given by
$G(x)=\frac{1+2x+2x^{2}+2x^{3}+2x^{4}+4x^{5}+3x^{6}+2x^{7}+x^{8}+x^{9}}{1-x^{2}-x^{3}-x^{4}+x^{5}-x^{6}-x^{7}-x^{8}-x^{10}}.$ (15)
Proof.
In this case, the $\mathcal{Q}$-comb is a $(2,3,1)$-comb and the metatile-generating digraph (Fig. 1) contains 4 inner cycles of which two are errant loops. Theorem 6 can therefore be applied and we choose the $10^{3}$ 1 node as the pseudo-common node $\mathcal{P}$. The inner cycles passing through $\mathcal{P}$ ($S^{2}C_{[4]}$ and $C_{[2]}SC_{[4]}$) have lengths $L_{1}=4$ and $L_{2}=6$ and their respective errant loops ($C_{[3]}$ and $C_{[2]}$) have lengths $L_{3}=3$ and $L_{4}=2$. The common circuits ($C_{[6]}S^{3}$ and $C_{[6]}C_{[2]}S^{2}$) have lengths $l_{\rm c1}=6$ and $l_{\rm c2}=8$. There is just one outer cycle ($S$) and its length $l_{\rm o1}=1$. From (4) we have
$B_{n}=\delta_{n,0}-\delta_{n,2}-\delta_{n,3}-\delta_{n,4}+\delta_{n,5}+\delta_{n,9}+B_{n-1}+B_{n-2}-2B_{n-5}+2B_{n-6}-B_{n-9}+B_{n-10}-B_{n-11}.$
The generating function for $S_{n}$ is therefore
$G(x)=\frac{1}{x^{5}}\biggl{(}\frac{1-x^{2}-x^{3}-x^{4}+x^{5}+x^{9}}{1-x-x^{2}+2x^{5}-2x^{6}+x^{9}-x^{10}+x^{11}}-\frac{1-x^{5}}{1-x}\biggr{)},$
which reduces to (15). ∎
The proof of the second bijection (Theorem 11) requires the following bijection established in [10] between strongly restricted permutations and tiling with $(\frac{1}{2},g)$-fences (which are just $(\frac{1}{2},g,\frac{1}{2})$-combs) where $g$ is always a nonnegative integer. As the teeth (originally called posts in the context of fences) are not of integer width, it is convenient to regard each cell as being divided into two equal slots into which a tooth can fit [11].
Lemma 10.
There is a bijection between (i) the permutations of $\mathbb{N}_{n}$ satisfying $\pi(i)-i\in\mathcal{D}$ for each $i\in\mathbb{N}_{n}$ and (ii) the tilings of an $n$-board using $(\frac{1}{2},d_{j})$-fences with the left tooth always placed in a left slot and $d_{j}$ equal to the non-negative elements of $\mathcal{D}$ and $(\frac{1}{2},-d_{k}-1)$-fences with their gaps aligned with cell boundaries and $d_{k}$ equal to the negative elements of $\mathcal{D}$. If the left slot of cell $i$ of the $n$-board is occupied by the left (right) tooth of a $(\frac{1}{2},g)$-fence then $\pi(i)=i+g$ ($\pi(i)=i-g-1$). For example, the tiling in Fig. 3 corresponds to the permutation
$\biggl{(}\!\!\begin{array}[]{ccccccccccccc}1&2&3&4&5&6&7&8&9&10&11&12&13\\ 1&11&3&4&2&6&7&5&9&10&8&12&13\end{array}\!\!\biggr{)}.$
We refer to a fence corresponding to an excedance (i.e., a position $i$ such that $\pi(i)>i$) as an up fence and a fence giving $\pi(i)-i<0$ as a down fence. The case of a fixed point ($\pi(i)=i$) corresponds to a gapless fence which is just an ordinary square tile (that must be aligned with a cell).
The following bijection is an extension of that noted but not proved by Baltić [7]. The proof we give requires the notion of a comb with $t$ teeth all of length $w$ separated by gaps of length $g$ which we refer to as a $(w,g;t)$-comb [5, 4].
Theorem 11.
There is a bijection between the $k$-subsets of $\mathbb{N}_{n}$ with disallowed differences $\mathcal{Q}=\{m,2m,\ldots,jm\}$, where $j,m\geq 1$, and the permutations $\pi$ of $\mathbb{N}_{n+jm}$ satisfying $\pi(i)-i\in\{-m,0,jm\}$ for all $i\in\mathbb{N}_{n+jm}$ that contain $k$ excedances.
Figure 3: A 13-board tiled with an up $(\frac{1}{2},9)$-fence ($F$), three down $(\frac{1}{2},2)$-fences ($\bar{F}$), and 9 squares.
Proof.
From Lemma 10, the number of such permutations is the number of ways to tile an $(n+jm)$-board using squares aligned with the cell boundaries ($S$), up $(\frac{1}{2},jm)$-fences ($F$), and down $(\frac{1}{2},m-1)$-fences ($\bar{F}$). Suppose the left tooth of the first $F$ on the board occupies the left slot of cell $i$. The only possible way to fill the right slot of cell $i$ is with the left tooth of an $\bar{F}$. The right tooth of the $\bar{F}$ lies on the left slot of cell $i+m$. If $j=1$, the right slot of this cell is occupied by the right tooth of the first $F$. Otherwise it must again be occupied by the left tooth of another $\bar{F}$ and similarly until the right tooth of the first $F$ is reached at cell $i+jm$. Hence this $F$ has $j$ $\bar{F}$ placed end-to-end in its interior. Combined, these $j+1$ tiles occupy exactly the same cells as a $(1,m-1;j+1)$-comb (Fig. 3 shows an example when $j=m=3$). All subsequent $F$ on the board behave similarly. Note that owing to the fact that any $\bar{F}$ must have its left tooth in the right slot of a cell, such tiles can only appear on the board in conjunction with $j-1$ other $\bar{F}$ surrounded by an $F$. Hence there is a bijection between tilings of an $(n+jm)$-board using $k$ $F$, $jk$ $\bar{F}$ and $n+jm-(j+1)k$ $S$ and the tilings of a board of the same length using $k$ $(1,m-1;j+1)$-combs and $n+jm-(j+1)k$ squares. ∎
Corollary 12.
Let $P^{\mathcal{D}}_{n}$ denote the number of permutations $\pi$ of $\mathbb{N}_{n}$ such that $\pi(i)-i\in\mathcal{D}$ for all $i\in\mathbb{N}$ and $P^{\mathcal{D}}_{n,k}$ denote the number of such permutations with $k$ excedances. Then for $m,j\geq 1$, $i,k\geq 0$, and $r=0,\ldots,m-1$, we have
$P^{\{-m,0,jm\}}_{im+r}$ $=\bigl{(}f^{(j+1)}_{i}\bigr{)}^{m-r}\bigl{(}f^{(j+1)}_{i+1}\bigr{)}^{r},$ (16)
$P^{\{-m,0,jm\}}_{im+r,k}$ $=[x^{k}]\biggl{(}\bigl{(}f^{(j+1)}_{i}(x)\bigr{)}^{m-r}\bigl{(}f^{(j+1)}_{i+1}(x)\bigr{)}^{r}\biggr{)},$ (17)
where $f^{(t)}_{n}$ is the $n$-th $(1,t)$-bonacci number defined by $f^{(t)}_{n}=f^{(t)}_{n-1}+f^{(t)}_{n-t}+\delta_{n,0}$, $f^{(t)}_{n<0}=0$, the $(1,t)$-bonacci polynomial $f^{(t)}_{n}(x)$ is defined by $f^{(t)}_{n}(x)=f^{(t)}_{n-1}(x)+xf^{(t)}_{n-t}(x)+\delta_{n,0}$, $f^{(t)}_{n<0}(x)=0$, and $[x^{k}]R(x)$ denotes the coefficient of $x^{k}$ in $R(x)$.
Proof.
Corollary 11 in [2], for $l\geq 0$, $m\geq 1$, $t\geq 2$, and $r=0,\ldots,m-1$, gives us
$S^{\{m,2m,\ldots,(t-1)m\}}_{lm+r}=\bigl{(}f^{(t)}_{l+t-1}\bigr{)}^{m-r}\bigl{(}f^{(t)}_{l+t}\bigr{)}^{r}.$ (18)
As a consequence of the bijection given in Theorem 11, $P^{\{-m,0,jm\}}_{n+jm}=S^{\{m,2m,\ldots,jm\}}_{n}$ and so $P^{\{-m,0,jm\}}_{n}=S^{\{m,2m,\ldots,jm\}}_{n-jm}$. Replacing $n$ by $im+r$ in this and using (18) with $t$ replaced by $j+1$ and $l$ replaced by $i-t+1$ gives (16). Similarly, Corollary 10 in [2] gives
$S^{\{m,2m,\ldots,(t-1)m\}}_{lm+r,k}=[x^{k}]\biggl{(}\bigl{(}f^{(t)}_{l+t-1}(x)\bigr{)}^{m-r}\bigl{(}f^{(t)}_{l+t}(x)\bigr{)}^{r}\biggr{)}.$
Combining this with the result from Theorem 11 that $P^{\{-m,0,jm\}}_{n+jm,k}=S^{\{m,2m,\ldots,jm\}}_{n,k}$ in the same way gives (17). ∎
Sequences in the OEIS that relate to Theorem 11 and Corollary 12 along with the corresponding $\mathcal{D}$ are as follows: A000045 $\{-1,0,1\}$, A006498 $\{-2,0,2\}$, A006500 $\{-3,0,3\}$, A031923 $\{-4,0,4\}$, A224809 $\{-2,0,4\}$, A224810 $\{-3,0,6\}$, A224811 $\{-2,0,8\}$, A224812 $\{-2,0,10\}$, A224813 $\{-2,0,12\}$, A224814 $\{-3,0,9\}$, A224815 $\{-4,0,8\}$.
4 Connections with compositions
A composition of a positive integer $n$ is a way of expressing $n$ as a sum of positive integers (referred to as parts) where the order of the parts is significant; thus, for example, 1+2 and 2+1 are different compositions of 3. For convenience, it is customary to take the number of compositions of 0 (any negative integer) as being 1 (0). It is well known that the number of compositions of $n$ is the same as the number of tilings of an $n$-board using $m$-ominoes for arbitrary nonnegative integers $m$ and hence that the number of compositions of $n$ into distinct parts $m_{1},m_{2},\ldots,m_{i},\ldots$ is given by $B_{n}=\delta_{n,0}+\sum_{i=1}B_{n-m_{i}}$, $B_{n<0}=0$ for all integers $n$. It follows that any tiling in which all possible metatiles have distinct lengths corresponds to a composition into parts equal to the lengths of the metatiles. We first examine two classes of $\mathcal{Q}$ where there are a finite number of possible metatiles when restricted-overlap tiling with squares and $\mathcal{Q}$-combs. Recall that the number of possible metatiles is finite iff $2r\geq q$, where $r$ is the length of the rightmost tooth of the $\mathcal{Q}$-comb (Lemma 2 in [1]). It is sometimes more convenient to express this condition as $r\geq s-1$, where $s=q+1-r$ is the sum of the lengths of the teeth and gaps up to the rightmost tooth.
Lemma 13.
Suppose $q>0$ is the largest element of $\mathcal{Q}$. Then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts which are drawn from a finite set of distinct parts iff every metatile, when restricted-overlap tiling with squares and $\mathcal{Q}$-combs, contains no more than two combs. Then the parts are 1, $q+1$, and $q+1+p_{i}$, where $p_{i}$ is the cell number of the $i$-th empty cell in the $\mathcal{Q}$-comb.
Proof.
If there are a finite number of parts, there are a finite number of metatiles which means that, in all cases, the start of the final comb in a metatile containing more than one comb must lie within the first comb (see the proof of Lemma 2 in [1]). If there were a third comb present, the cells it occupies without overlapping (of which there must be at least one, i.e., its cell 0) could be replaced by squares and give rise to a metatile of the same length. The parts would then not be distinct. The square and filled comb metatiles are of length 1 and $q+1$, respectively. The remaining metatiles contain two combs. If the start of the second comb is in cell $p_{i}$ of the first comb, the resulting metatile (completed by filling any remaining empty cells with squares) is easily seen to have a length of $q+1+p_{i}$. ∎
In the proofs of the following two theorems we employ combs that, in general, have a periodic pattern of teeth and gaps before the rightmost tooth is reached. Extending the notation used for combs with teeth all the same length and gaps all the same length [4, 5], we call an $(l,g,l,g,\ldots,r)$-comb with $t$ teeth an $(l,g,r;t)$-comb and we refer to an $(l,g,m,h,l,g,m,h,\ldots,r)$-comb with $t$ teeth as an $(l,g,m,h,r;t)$-comb. Note that an $r$-omino (an $(l,g,r)$-comb) can be regarded as an $(l,g,r;t)$- or $(l,g,m,h,r;t)$-comb with $t=1$ ($t=2$).
Theorem 14.
If
$\mathcal{Q}=\{0,1,\ldots,l-1,l+g,l+g+1,\ldots,2l+g-1,2(l+g),2(l+g)+1,\ldots,\\ (t-1)(l+g),\ldots,(t-1)(l+g)+r-1=q>0\}$
for some $l\geq g>0$, $t\in\mathbb{Z}^{+}$, and, for $t>1$, $r\geq(t-1)(l+g)-1$, then $S_{n}^{\mathcal{Q}}$ is the number of compositions of $n+q$ into parts 1 and $q+1$ when $t=1$ and parts 1, $q+1$, $q+l+1,\ldots,q+l+g$, $q+2l+g+1,\ldots,q+2(l+g),\ldots,q+(t-1)l+(t-2)g+1,\ldots,q+(t-1)(l+g)$ otherwise. Note that $\mathcal{Q}$ reduces to $\{0,1,\ldots,r-1=q\}$ when $t=1$.
Figure 4: Construction of instances of metatiles containing two combs when the combs are (a) (2,2,7;3)-combs, (b) (5,2,1,2,9;3)-combs, and (c,d) (2,2,3,2,8;3)-combs. In each case, the second comb is displaced downwards a little from its final position so that parts of the cells of the first comb it overlaps are visible.
Proof.
The $\mathcal{Q}$-comb in this case is an $(l,g,r;t)$-comb. When $t=1$, there are no two-comb metatiles; the condition $r>1$ ensures that the filled comb metatile (which in this case is just the comb) is not the same length as the square. For $t>1$, the condition $r\geq(t-1)(l+g)-1$ is the same as $2r\geq q$ and thus establishes that there are a finite number of metatiles. Then the condition $l\geq g$ ensures that a metatile can have at most two combs (Fig. 4(a)). This is because if $g>l$, the leftmost tooth of a comb and the leftmost cell of another comb could both be placed within the first gap of the initial comb. The result then follows from Lemma 13. ∎
OEIS sequences corresponding to a $\mathcal{Q}$ that Theorem 14 applies to along with the set itself are as follows: A00045 {1}, A006498 {2}, A000930 {1,2}, A079972 {2,3}, A003269 {1,2,3}, A351874 {1,3,4}, A121832 {2,3,4}, A003520 {1,2,3,4}, A259278 {2,3,4,5}, A005708 {1,2,3,4,5}, A276106 {2,3,4,5,6}, A005709 {1,2,3,4,5,6}, A322405 {2,3,4,5,6,7}, A005710 {1,2,3,4,5,6,7}, A368244 {2,3,4,5,6,7,8}.
Theorem 14 covers all cases where there are a finite number of metatiles for $q\leq 7$. The following generalization of the theorem (in which the $\mathcal{Q}$-combs are periodic in the first two teeth and gaps up to the final tooth) extends this to $q\leq 11$.
Theorem 15.
If
$\mathcal{Q}$ =$=\{0,1,\ldots,l-1,l+g,l+g+1,\ldots,l+g+m-1,$
$\qquad l+g+m+h,l+g+m+h+1,\ldots,2l+g+m+h-1,$
$\qquad 2(l+g)+m+h,2(l+g)+m+h+1,\ldots,2(l+g+m)+h-1,\ldots,*\},$
* =$=\begin{cases}\dfrac{t-1}{2}(l+g+m+h),\ldots,\dfrac{t-1}{2}(l+g+m+h)+r-1=q>0,&\text{t odd};\\ \dfrac{t}{2}(l+g)+\Bigl{(}\dfrac{t}{2}-1\Bigr{)}(m+h),\ldots,\dfrac{t}{2}(l+g)+\Bigl{(}\dfrac{t}{2}-1\Bigr{)}(m+h)+r-1=q,&\text{t even};\end{cases}$
where $g,m,h,t\in\mathbb{Z}^{+}$, and, for $t>1$, $r\geq\frac{1}{2}(t-1)(l+g+m+h)-1$ if $t$ is odd and $r\geq\frac{1}{2}t(l+g)+(\frac{1}{2}t-1)(m+h)-1$ if $t$ is even, and either (i) $l\geq g+m+h$ or (ii) $l\geq g,h$ and $m\geq l-1+g$, then $S_{n}^{\mathcal{Q}}$ is the number of compositions of $n+q$ into parts 1 and $q+1$ when $t=1$ and, when $t>1$, parts 1, $q+1$, $q+l+1,\ldots,q+l+g$, $q+l+g+m+1,\ldots,q+l+g+m+h,\ldots,*$, where
={$*=\begin{cases}q+\dfrac{t-1}{2}(l+g+m)+\dfrac{t-3}{2}h+1,\ldots,q+\dfrac{t-1}{2}(l+g+m+h),&\text{t odd};\\ q+\dfrac{t}{2}l+\Bigl{(}\dfrac{t}{2}-1\Bigr{)}(g+m+h)+1,\ldots,q+\dfrac{t}{2}(l+g)+\Bigl{(}\dfrac{t}{2}-1\Bigr{)}(m+h),&\text{t even}.\end{cases}$
Proof.
Here the $\mathcal{Q}$-comb is an $(l,g,m,h,r;t)$-comb. The $t=1,2$ cases are covered by the proof of Theorem 14. First note that if $t$ is odd, the length of the comb $q+1=\frac{1}{2}(t-1)(l+g+m+h)+r$ and it is $\frac{1}{2}t(l+g)+(\frac{1}{2}t-1)(m+h)+r$ if $t$ is even from which the respective conditions on $r$ (so that there are a finite number of metatiles) follow. There are two mutually exclusive sets of conditions that ensure that the metatiles contain no more than two combs. Condition (i) that $l\geq g+m+h$ (see Fig. 4(b)) arises as a result of the observation that if $l$ were one less than $g+m+h$ and a second comb were placed starting at the first gap of the first comb, the start a third comb could be placed starting at the final cell of the second gap of the first comb. The first part of condition (ii), namely that $l\geq g,h$ is for the same reasons as the $l\geq g$ condition in Theorem 14. We must also have $l+g+m\geq g+m+h$ (see Fig. 4(c)) which also implies that $l\geq h$. Were $l+g+m$ one less than $g+m+h$, it can be seen that after placing the second comb at the start of the first gap, a third could be placed at the end of the second gap of the first comb. At the same time, we must ensure that $m$ is not too small; otherwise the start of a third comb could be inserted at the start of the second gap of the first comb. Thus $l-1+g\leq m$ (see Fig. 4(d)). The result then follows from Lemma 13. ∎
We now turn to compositions where the parts are drawn from an infinite set. In the context of tiling this means that there is an infinite number of possible metatiles and so the metatile-generating digraph has an inner cycle.
Lemma 16.
If the metatile-generating digraph for a tiling contains a common node and more than one inner cycle then there exist distinct metatiles of the same length.
Proof.
Suppose the digraph contains inner cycles $I_{1}$ and $I_{2}$. If the sequence of nodes and arcs to create one metatile involves executing $I_{1}$ followed by $I_{2}$ then a distinct metatile of the same length can be generated by swapping the order in which the inner cycles are traversed. ∎
Four classes of $\mathcal{Q}$ whose associated digraphs possess common nodes have been investigated (Theorems 5–8 in [1]). One of these classes (Theorem 8 in [1]) always has at least two inner cycles; the simplest member of this class is $\mathcal{Q}=\{1,4\}$ (see Fig. 2). Another class (Theorem 7 in [1]) has just one inner cycle but always has the metatiles $C_{[q+1]}C_{[l]}^{2}$ and $C_{[q+1]}S^{a-1}C_{[2l]}S$, where $a=\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert$, which are of the same length (see Fig. 5 in [1]). The metatile-generating digraph of one of the remaining classes has $a$ inner cycles (Theorem 5 in [1]). The $a=1$ case leads to the following result.
Theorem 17.
If $\mathcal{Q}=\{1,\ldots,p-1,p+1,\ldots,q\}$, where $2p>q$, then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts 1 and $q+1+jp$ for $j=0,1,2,\ldots$.
Figure 5: Digraph for restricted-overlap tiling with squares ($S$) and $(p,1,q-p)$-combs ($C$), where $2p>q$. This corresponds to the $\mathcal{Q}=\{1,\ldots,p-1,p+1,\ldots,q\}$ case.
Proof.
From the digraph (Fig. 5) it can be seen that the metatiles are $S$ and $C_{[q+1]}C_{[p]}^{j}S$ for $j=0,1,2,\ldots$ which are of lengths 1 and $q+1+jp$, respectively. ∎
OEIS sequences corresponding to a $\mathcal{Q}$ that Theorem 17 applies to along with the set itself are as follows: A130137 {1,3}, A317669 {1,2,4}, A375185 {1,2,3,5}, A375186 {1,2,4,5}.
The following theorem originates from Theorem 6 in [1]. Although there appears to be only one sequence in the OEIS that it applies to (namely, A224809 for which $\mathcal{Q}$ is {2,4}), there are 18 other cases of $\mathcal{Q}$ it applies to for $q\leq 9$ [1].
Theorem 18.
Let $\theta$ be the bit string representation of $\mathcal{Q}$ whereby the $j$-th bit from the right of $\theta$ is 1 if and only if $j\in\mathcal{Q}$. By $\lfloor\theta/2^{b}\rfloor$ we mean discarding the rightmost $b$ bits in $\theta$ and shifting the remaining bits to the right $b$ places. Let $p_{i}\in\mathbb{N}_{q}-\mathcal{Q}$ for $i=1,\ldots,a$ and $p_{i}<p_{i+1}$. Using $\mid$ to denote the bitwise OR operation, if $\theta\mid\lfloor\theta/2^{p_{i}-1}\rfloor$ for each $i=1,\ldots,a-1$ is all ones after discarding the leading zeros, $a\geq 2$, and $p_{a}=q-r$ (which implies that $r\geq 1$), then if (a) $q=2r+1$ or (b) $q>2r+1$ and $1\leq p_{a-1}\leq r$, then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts 1, $q+1$, $q+1+p_{i}$ for $i=1,\ldots,a-1$, and $q+1+j(q-r)$ for $j\in\mathbb{Z}^{+}$.
Figure 6: Digraph for tiling a board with squares and combs corresponding to $\mathcal{Q}$ specified in Theorem 18. $\sigma_{i}$ is the bit string corresponding to filling the first $i-1$ empty cells of a comb with squares and discarding the leading 1s.
Proof.
The digraph (Fig. 6, [1]) generates the metatiles $S$, $C_{[q+1]}S^{a}$, $C_{[q+1]}S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a-1$, and $C_{[q+1]}S^{a-1}C_{[q-r]}^{j}S$ for $j\in\mathbb{Z}^{+}$. Their respective lengths are 1, $q+1$, $q+1+p_{i}$, and $q+1+j(q-r)$. As $q-r=p_{a}>p_{a-1}$, the lengths are all distinct. ∎
5 Connections with bit strings
Others have noted that the number of length-$n$ bit strings (or binary words) having no two 1s that have positions in the string that differ by $q$ (or, equivalently, are separated from one another by $q-1$ digits) is $S^{\{q\}}_{n}$ [15]. This can be generalized by considering the following bit string $s$ representing subset $\mathcal{S}$: the $j$-th bit from the right of $s$ is 1 if and only if $j\in\mathcal{S}$. (This representation is used in the program in the Appendix.) Then it is clear that $S_{n,k}$ is the number of length-$n$ bit strings that contain $k$ 1s placed in such a way that the difference in positions of any two 1s does not equal an element of $\mathcal{Q}$. We now mention an instance of this that corresponds to A130137 in the OEIS [20]. If one instead places the same restrictions on 0s rather than 1s, for the $\mathcal{Q}=\{1\}$ case, the bit strings (i.e., those containing no 00) are called Fibonacci binary words, so named as the number of such words of length $n$ is $F_{n+2}$. A130137 is described as the number of length-$n$ Fibonacci binary words that do not contain the substring 0110. This is $S^{\{1,3\}}_{n}$ since the three other possible disallowed substrings that only prevent the occurrence of two 0s whose positions differ by 3 (i.e., $0$, 0010, and 0100) do not need to be mentioned; they each contain 00 and so cannot be present in Fibonacci binary words.
The final bijection we present concerns equivalence classes for the occurrence of a subword within binary words. We say that two binary words of the same length $n$ are equivalent with respect to a subword $\omega$ if $\omega$ occurs in the same position(s) in those words. One can represent the equivalence class by the set of subword positions, counting the leftmost position as 1. For example, all binary words of the form $xx10010010xx$, where each $x$ can be 0 or 1, belong to the same equivalence class with respect to the subword $10010$ and this class is represented by the set $\{3,6\}$.
Theorem 19.
Let $\omega$ be a length-$l$ binary subword. We construct set $\mathcal{Q}$ as follows: $j\in\mathcal{Q}$ iff $\lfloor\omega/2^{j}\rfloor\neq\omega\bmod 2^{l-j}$ for $j=1,\ldots,l-1$. If $q$ is the largest element of $\mathcal{Q}$ (and taking $q$ as zero when $\mathcal{Q}=\{\}$) and $q=l-1$, then the number of equivalence classes of binary words of length $n$ with respect to subword $\omega$ is $S^{\mathcal{Q}}_{n-q}$. Furthermore, $S^{\mathcal{Q}}_{n-q,k}$ is the number of such equivalence classes whose representation as a set is of size $k$.
Proof.
We show that the $k$-subsets of $\mathbb{N}_{n-q}$ satisfying the conditions for disallowed differences specified by $\mathcal{Q}$ are the same as the sets representing the equivalence classes of binary words of length $n$ in which $\omega$ appears $k$ times. In general, the possible positions of any subword of length $l$ in a word of length $n$ (and therefore the possible elements of the set representing an equivalence class) are $1,\ldots,n-l+1=n-q$. From a restricted subset $\mathcal{S}$ of $\mathbb{N}_{n-q}$ we construct an equivalence class of words as follows. Starting with a length-$n$ word of $x$s (each $x$ representing 0 or 1), for each $s\in\mathcal{S}$ we place the start of subword $\omega$ at position $s$ in the word. Note that $\lfloor\omega/2^{j}\rfloor$ gives the $l-j$ leftmost bits of $\omega$ and $\omega\bmod 2^{l-j}$ gives the $l-j$ rightmost bits. If the two are equal (and so $j\notin\mathcal{Q}$) then, on shifting the digits of subword $\omega$ by $j$ places to the right and overlaying them on the original subword, none of the digits of the original subword are changed. This means that we have constructed a valid equivalence class once we have specified the restrictions on the remaining $x$s to avoid any further appearances of $\omega$. Note that if there an $\omega$ at position $i$ and no other at $i+j$ for all $j=1,\ldots,q$, we may place another $\omega$ at position $i+l$ or after this. This coincides with the fact that $l=q+1$ or any larger integers are not contained in the set $\mathcal{Q}$ of disallowed shifts. The reversible nature of the construction establishes the bijection. ∎
[…]
In the rest of this section we show that the elements of $\mathcal{Q}$ concerned with the bijection given in Theorem 19 are always a well-based sequence and hence the generating function for the number of equivalence classes and the recursion relation for the number of equivalence classes whose set representation is of size $k$ are straightforward to obtain. The sequence $q_{1},q_{2},\ldots,q_{m}$, where $q_{i}<q_{i+1}$ for all $i$, is said to be a well-based sequence if $q_{1}=1$ and for all $j>1$ and $\Delta=1,\ldots,q_{j}-1$, there is some $q_{i}$ such that $q_{j}=q_{i}+\Delta$ [13, 21]. In the present context, we can regard $q_{i}$ as being the elements of $\mathcal{Q}$ with $q_{m}=q$, its largest element. In the proof of Theorem 20, we will use the following definition which is equivalent aside from also classing the empty set as being well based: the sequence is well based if $a=\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert$ is zero or if, for all $i,j=1,\ldots,a$ (where $i$ and $j$ can be equal), $p_{i}+p_{j}\notin\mathcal{Q}$, where the $p_{i}$ are the elements of $\mathbb{N}_{q}-\mathcal{Q}$. E.g., the only well-based sequences of length 3 are the elements of the sets $\{1,2,3\}$, $\{1,2,4\}$, $\{1,2,5\}$, and $\{1,3,5\}$.
Theorem 20.
If $\mathcal{Q}$ is constructed from subword $\omega$ as described in Theorem 19 and $q$ is the largest element of $\mathcal{Q}$, then the elements of $\mathcal{Q}$ are a well-based sequence with the additional more general property that $\Lambda\notin\mathcal{Q}$ where $\Lambda$ is any linear combination (with integer coefficients) of the $p_{i}$.
Proof.
Suppose $\omega$ is the length-$l$ bit string $b_{l}b_{l-1}\cdots b_{2}b_{1}$. Then $p_{i},p_{j}\notin\mathcal{Q}$ iff $b_{n}=b_{n-p_{i}}$ and $b_{n}=b_{n-p_{j}}$ for $n\leq l$ with $n\geq p_{i}+1$ and $n\geq p_{j}+1$, respectively. Hence we also have $b_{n}=b_{n-p_{i}-p_{j}}$ for $1+p_{i}+p_{j}\leq n\leq l$ which implies that $p_{i}+p_{j}\notin\mathcal{Q}$ and hence the elements of $\mathcal{Q}$ form a well-based sequence. In an analogous way, it is easily seen that we have $b_{n}=b_{n-\Lambda}$ where $\Lambda$ is any integer-coefficient linear combination of the $p_{i}$ such that $0<\Lambda<n$ and hence $\Lambda\notin\mathcal{Q}$. ∎
The additional property in Theorem 20 means, for example, that the sets $\{1,2,5\}$ (for which $p_{1}=3$, $p_{2}=4$) and $\{1,2,3,5,7\}$ (for which $p_{1}=4$, $p_{2}=6$) do not correspond to any $\omega$ (since in both cases $p_{2}-p_{1}\in\mathcal{Q}$) although their elements form well-based sequences.
To obtain Corollary 22 to Theorem 20, we require the following result.
Lemma 21 (Corollary 1 in [1]).
If the elements $q_{i}$ of $\mathcal{Q}$ are a well-based sequence then the generating function for $S^{\mathcal{Q}}_{n}$ is given by
$G(x)=\frac{c}{(1-x)c-x},$ (19)
where $c=1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}$.
Corollary 22.
The generating function $\tilde{G}(x)$ for $E^{(\omega)}_{n}$, the number of equivalence classes of length-$n$ binary words with respect to a length-$l$ subword $\omega$ is given by
$\tilde{G}(x)=\frac{\displaystyle{1-\sum_{i=1}^{\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert}\!\!x^{p_{i}}}}{\displaystyle{(1-x)\biggl{(}1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}\biggr{)}-x}},$ (20)
where $\mathcal{Q}$ is obtained from $\omega$ using the procedure in Theorem 19, $q_{i}$ are the elements of $\mathcal{Q}$ of which the largest is $q$ (with $q=0$ when $\mathcal{Q}=\{\}$), and $p_{i}$ are the elements of $\mathbb{N}_{q}-\mathcal{Q}$, provided that $q=l-1$.
Proof.
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and the sums in (20) are omitted leaving $\tilde{G}(x)=1/(1-2x)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n}=2^{n}$ which coincides with the values given by $\tilde{G}(x)$. For $q>0$, since $E^{(\omega)}_{n}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 19, $E^{(\omega)}_{n}=S^{\mathcal{Q}}_{n-q}$ for $n\geq q$, we have, on using Lemma 21, that
$\tilde{G}(x)=\frac{x^{q}c}{(1-x)c-x}+1+x+\cdots+x^{q-1}=\frac{x^{q}c}{(1-x)c-x}+\frac{1-x^{q}}{1-x}.$
This leads to (20) on using the result that $c+\sum_{i=1}^{\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert}x^{p_{i}}=x+x^{2}+\cdots+x^{q}$. ∎
Theorem 5 of [1] also gives a recursion relation for $B^{\mathcal{Q}}_{n,k}$ when the elements of $\mathcal{Q}$ form a well-based sequence. Using Theorems 19 and 1 it is therefore straightforward to obtain a recursion relation for the number of equivalence classes with respect to length-$l$ subword $\omega$ with associated set $\mathcal{Q}$ (satisfying $q=l-1$) of length-$n$ binary words that have a set representation of size $k$.
6 Discussion
Figure 7: Digraph for restricted-overlap tiling with squares ($S$) and $(2,4,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,6\}$ case.
Theorem 6, which we derived here in order to obtain the recursion relation for $B^{\{1,5\}}_{n}$, will also be used elsewhere to obtain recursion relations for $B^{\mathcal{Q}}_{n}$ and $B^{\mathcal{Q}}_{n,k}$ for four other classes of $\mathcal{Q}$; the next instance of $\mathcal{Q}$ it applies to is $\mathcal{Q}=\{1,2,5,7\}$. The metatile-generating digraph corresponding to the next case in the the $\{1,q\}$ series (Fig. 7) lacks the symmetry of the $\{1,5\}$ case and does not fall into any of the classes of digraph that we have been able to derive a recursion relation for so far. Unlike with strongly restricted permutations with the sole conditions of $\pi(i)-i\in\mathcal{D}$, where there is a general procedure for obtaining the generating function for the number of permutations [6], the literature does not appear to contain any analogous procedure for obtaining numbers of restricted combinations.
Appendix: C program for finding $S_{n}$ and $S_{n,k}$
There are two main features that contribute to the efficiency of the algorithm used in the program listed below. Rather than directly computing whether the difference of each pair $x,y$ of elements of subset $\mathcal{S}$ is in $\mathcal{Q}$, we perform a bitwise AND operation on bit string representations of $\mathcal{S}$ shifted $x$ places and $\mathcal{Q}$. If the AND operation gives an answer of 0 then $y-x$ does not equal any element of $\mathcal{Q}$ for all $y>x$. We also exploit the result that if $\mathcal{S}$ is an allowed subset of $\mathbb{N}_{n}$ then it is an allowed subset of $\mathbb{N}_{m}$ for any $m>n$.
// anything after // on the same line is a comment
// program name: rcl.c (restricted combinations on a line)
// purpose: counts subsets of {1,2,...,n} such that
// no two elements have a difference equal to an element of Q
[…]
// to create an executable called rcl using the GNU C compiler enter
// gcc rcl.c -o rcl
// include libraries
#include <stdio.h>
#include <stdlib.h>
#include <string.h>
// sizeof( ) returns the number of bytes used by a variable type
// the variable type unsigned int typically uses 4 bytes
// hence MAXn is typically 32
#define MAXn 8*sizeof(unsigned int)
// after including the libraries and defining MAXn, program starts here
// argc equals the number of command-line arguments plus 1
// argv is an array of character strings containing those arguments
int main(int argc,char **argv) {
// start by defining the variables used by the program
// j-th bit from the right in the bit string is 1 iff j is in the set it represents
unsigned int Q=0,s,ss; // bit strings giving Q, subset, and shifted subset
char *p; // used for reading in Q (given as a comma-separated list) from command line
[…]
i,j,k=0,test,lastn=0;
[…]
// end of definitions of variables
// give examples of usage if program run with no arguments
if (argc==1) { // rcl called with no arguments
[…]
exit(1); // end the program
} // end if
// p starts by pointing to the character before the start of the 1st argument
for (p=argv[1]-1;p;p=strchr(p,’,’)) { // loop to read in Q from command line argument
// line(s) within these braces only executed if p is not zero
// ++p means add 1 to p first
// atoi( ) converts a string into an integer, stopping as soon as it reaches a non-digit
// x<<m means shift x in binary m places to the left
// | is the bitwise OR operation
Q=Q|1<<(atoi(++p)-1); // add to existing Q
} // end of loop: p is updated to point to the next comma in argument or zeroed if none
// && means a logical AND
tri=(argc>2 && argv[2][0]==’t’); // true if argument 2 in command line is t
[…]
for (i=0;i<=MAXn;i++) {
// cond?x:y returns x if cond is true and y otherwise
for (j=0;j<=(tri?i+1:1);j++) {
S[i][j]=j?0:1; // set k=0 totals to 1, zero k>0 totals
// ~0 means all ones in binary
for (s=1;s<~0;s++) { // loop to test all possible nonempty subsets s
if (tri) {
[…]
for (k=0,ss=s;ss;ss>>=1) {
if (ss&1) k=k+1; // & is the bitwise AND operation
for (ss=s,i=1;i<=MAXn;i++) { // test subset s
test=ss&1; // will only test shifted subset against Q if rightmost bit of ss is 1
ss>>=1; // shift shifted subset ss to the right by 1
if (ss==0) { // s is ok (no disallowed differences found)
// i is now minimum n for which s is ok
if (i>lastn) { // total(s) for previous n finished
for (j=0;S[lastn][j];j++) {
[…]
if (tri) puts(""); // new line
else fflush(stdout); // display immediately
lastn=i; // update lastn for next time it is used
for (j=i;j<=MAXn;j++) {
S[j][k]++; // if s is ok for a given n, it is ok for all higher n
break; // done with current s
if (test && (Q&ss)) break; // disallowed difference found - done with current s
for (j=0;S[lastn][j];j++) {
printf("%d,",S[lastn][j]); // lastn now equals MAXn
puts(""); // print a new line at the end